# PiLocNet: Physics-informed neural network on 3D localization with rotating point spread function

### v1

@@ Abstract
For the 3D localization problem using point spread function (PSF) engineering, we propose a novel enhancement of our previously introduced localization neural network, LocNet. The improved network is a physics-informed neural network (PINN) that we call PiLocNet. Previous works on the localization problem may be categorized separately into model-based optimization and neural network approaches. Our PiLocNet combines the unique strengths of both approaches by incorporating forward-model-based information into the network via a data-fitting loss term that constrains the neural network to yield results that are physically sensible. We additionally incorporate certain regularization terms from the variational method, which further improves the robustness of the network in the presence of image noise, as we show for the Poisson and Gaussian noise models. This framework accords interpretability to the neural network, and the results we obtain show its superiority. Although the paper focuses on the use of single-lobe rotating PSF to encode the full 3D source location, we expect the method to be widely applicable to other PSFs and imaging problems that are constrained by known forward processes.
keywords:
3D localization, point spread function engineering, physics-informed neural network, inverse problems
[…]
@@ 1 Introduction
Inspired by the idea of PINN, we propose here the Physics-Informed Localization Network (PiLocNet) method for the PSF localization problem. Given the process of PSF image generation, the forward model is a known piece of physical information. Supplying such physical information to the neural network is helpful to the network in generating better results than a random black-box kind of fitting approach characteristic of more conventional neural networks. With model-specific data fitting and regularization terms, a PINN-based method is interpretable to the neural network. The numerical experiments we report here have been conducted based on the RPSF model. Our results, as we will show, prove that the added physical information can significantly improve the prediction accuracy in terms of both precision and recall rates. Our ablation studies also verified the robustness of PiLocNet.
@@ 2.1 Forward model of single-lobe RPSF
The forward model has been described in great detail in our previous papers [9, 13, 14] to which we refer the reader. Here, we only present a brief summary of the model. The RPSF image, $\mathcal{A}_{\zeta}$, for a point source with defocus parameter $\zeta$, a unit flux $f=1$, at the source location $\mathbf{r}_{O}=(x_{O},y_{O})$ is given by:
@@
where $\zeta=-\frac{\pi\delta zR^{2}}{\lambda z_{O}(z_{O}+\delta z)}$, $\Omega$ represents the circular disk-shaped clear pupil and $i=\sqrt{-1}$. The quantity, $\mathbf{s}=\frac{\mathbf{r}}{\lambda z_{I}/R}$, is the position vector, $\mathbf{r}$, of an image-plane point relative to the Gaussian image location, when expressed in units of the Rayleigh diffraction scale, $\lambda z_{I}/R$; $\lambda$ is the imaging wavelength; $R$ is the radius of the pupil; $\delta z$, $z_{O}$, $z_{I}$ are the distances from the object plane to the in-focus object plane, the in-focus object plane to the pupil plane, and the pupil plane to the image plane, respectively. The symbol $\mathbf{u}$ denotes the position vector in the plane of the pupil, in units of the radius of the pupil. Its polar coordinates are $\mathbf{u}=(u,\phi_{u})$. The circular pupil is segmented into $L$ different contiguous annular Fresnel zones, with each zone carrying a spiral phase function, $\psi(\mathbf{u})$, with the number of complete phase cycles changing successively by 1 from one zone to the next. The RPSF can be shown [9] to continuously rotate within the scaled defocus range, $\zeta\in[-\pi L,\pi L]$, as it begins to spread out, break apart, and lose its shape unacceptably outside this range.
With the above formulation, for $N$ point sources the observed image count $G(x,y)$ at location $(x,y)$ may be expressed as:
@@
where $(x_{i},y_{i})$ and $f_{i}$ are the transverse coordinates and flux of the $i$th point source. Its depth coordinate, $z_{i}$, is embedded in the PSF $A_{i}$ and $\mathcal{N}$ is the operator for incorporating noise.
@@ 3.2 The pipeline of PiLocNet and its architecture
LocNet is a data-driven approach that only considered the Poisson noise scenario in Ref. [15]. Here we incorporate PINN into LocNet and propose a new framework, PiLocNet, for both Poisson and Gaussian noise cases. The main idea of PINN is to incorporate information about the physics of the problem at hand into the neural network’s loss function. This approach helps the training process to achieve accurate results by minimizing the loss with improved guidance provided by known physical information. In this context, we discuss how this concept can be implemented to solve the PSF problem. We modify the LocNet loss function by adding to it two extra terms,
@@
For most of the experiments conducted in the paper, the weights $w_{1},w_{2},w_{3}$, was in the ratio 1:700:1000 for the case (5) of Gaussian noise and 1:1:500 for the case (6) of Poisson noise. For experiments in ablation study on different noise levels, one can go through a searching process to reach optimal weight values correspondingly.
@@ 5 Conclusions
5.1 Disclosures
The authors declare no conflicts of interest.
5.2 Code, Data, and Materials Availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
5.3 Acknowledgments
Funding information: HKRGC Grants Nos. N_CityU214/19, CityU11301120, C1013-21GF, and CityU11309922; the Natural Science Foundation of China No. 12201286; Guangdong Basic and Applied Research Foundation 2024A1515012347; Shenzhen Science and Technology Program 20231115165836001; CityU Grant 9380101.

### v2

@@ Abstract
For the 3D localization problem using point spread function (PSF) engineering, we propose a novel enhancement of our previously introduced localization neural network, LocNet. The improved network is a physics-informed neural network (PINN) that we call PiLocNet. Previous works on the localization problem may be categorized separately into model-based optimization and neural network approaches. Our PiLocNet combines the unique strengths of both approaches by incorporating forward-model-based information into the network via a data-fitting loss term that constrains the neural network to yield results that are physically sensible. We additionally incorporate certain regularization terms from the variational method, which further improves the robustness of the network in the presence of image noise, as we show for the Poisson and Gaussian noise models. This framework accords interpretability to the neural network, and the results we obtain show its superiority. Although the paper focuses on the use of a single-lobe rotating PSF to encode the full 3D source location, we expect the method to be widely applicable to other PSFs and imaging problems that are constrained by well modeled forward processes.
journal: opticajournalarticletype: Research Article
@@ 1 Introduction
Figure 1: Overview of the PiLocNet: the main improvement is in the inclusion of a forward loss term based on the matrix $\mathcal{A}$ that models the 3D point spread function along with an appropriate regularization term into the loss function. This added physics information improves network training.
@@
Inspired by the idea of PINN, we propose here the Physics-Informed Localization Network (PiLocNet) method for the PSF localization problem, as shown schematically in Fig. 1. Given the process of PSF image generation, the forward model is a known piece of physical information. Supplying such physical information to the neural network is helpful to the network in generating better results than a random black-box kind of fitting approach characteristic of more conventional neural networks. With model-specific data fitting and regularization terms, a PINN-based method is interpretable to the neural network. The numerical experiments we report here have been conducted based on the RPSF model. Our results, as we will show, prove that the added physical information can significantly improve the prediction accuracy in terms of both precision and recall rates. Our ablation studies also verified the robustness of PiLocNet.
@@ 2.1 Forward model of single-lobe RPSF
The forward model has been described in great detail in our previous papers [9, 13, 14]. Here, we only present a brief summary of the model. The RPSF image, $\mathcal{A}_{\zeta}$, for a point source with defocus parameter $\zeta$, a unit flux $f=1$, at the source location $\mathbf{r}_{O}=(x_{O},y_{O})$ is given by:
@@
where $\zeta=-\frac{\pi\delta zR^{2}}{\lambda z_{O}(z_{O}+\delta z)}$, $\Omega$ represents the circular disk-shaped clear pupil and $i=\sqrt{-1}$. The quantity, $\mathbf{s}=\frac{\mathbf{r}}{\lambda z_{I}/R}$, is the position vector, $\mathbf{r}$, of an image-plane point relative to the Gaussian image location, when expressed in units of the Rayleigh diffraction scale, $\lambda z_{I}/R$, in which $\lambda$ is the imaging wavelength, $R$ is the radius of the pupil, and $\delta z$, $z_{O}$, $z_{I}$ are the distances from the object plane to the in-focus object plane, the in-focus object plane to the pupil plane, and the pupil plane to the image plane, respectively. The symbol $\mathbf{u}$ denotes the position vector in the plane of the pupil, in units of the radius of the pupil. Its polar coordinates are $\mathbf{u}=(u,\phi_{u})$. The circular pupil is segmented into $L$ different contiguous annular Fresnel zones, with each zone carrying a spiral phase function, $\psi(\mathbf{u})$, with the number of complete phase cycles changing successively by 1 from one zone to the next. The RPSF can be shown [9] to continuously rotate within the scaled defocus range, $\zeta\in[-\pi L,\pi L]$, as it begins to spread out, break apart, and lose its shape unacceptably outside this range. An illustration of the images of a single point source at different values of the depth parameter $\zeta$, when the RPSF is used, is presented in Fig  2.
Figure 2: single-lobe point spread function applied images with different depth $\zeta$.
With the above formulation, for $N$ point sources the observed image data count $G(x,y)$ at location $(x,y)$ may be expressed as:
@@
where $(x_{i},y_{i})$ and $f_{i}$ are the transverse coordinates and flux of the $i$th point source. Its depth coordinate, $z_{i}$, is embedded, via the depth-parameter value $\zeta_{i}$, in the PSF $A_{i}$ and $\mathcal{N}$ is the operator for incorporating noise.
@@ 3.2 The pipeline of PiLocNet and its architecture
LocNet is a data-driven approach that only considered the Poisson noise scenario in Ref. [15]. Here we incorporate PINN into LocNet and propose a new framework, PiLocNet, for both Poisson and Gaussian noise cases. The main idea of PINN is to incorporate information about the physics of the problem at hand into the neural network’s loss function, as illustrated in the Fig. 1. This approach helps the training process to achieve accurate results by minimizing the loss with improved guidance provided by known physical information. In this context, we discuss how this concept can be implemented to solve the PSF problem. We modify the LocNet loss function by adding to it two extra terms,
@@
For most of the experiments conducted in the paper, the weights $w_{1},w_{2},w_{3}$, were in the ratio 1:700:1000 for the case (5) of Gaussian noise and 1:1:500 for the case (6) of Poisson noise. For experiments in ablation study on different noise levels, one can go through a searching process to reach optimal weight values correspondingly.
@@ 5 Conclusions
[…]